OpenStudy (anonymous):

11x^2-18x=0 Please help

OpenStudy (shamim):

do u need solution of ur given equation

OpenStudy (shamim):

\[11x ^{2}-18x=0\]

OpenStudy (shamim):

\[or, x(11x-18)=0\]

OpenStudy (shamim):

\[or, x=0\]

OpenStudy (shamim):

\[or, 11x-18=0\]

OpenStudy (shamim):

\[or, x=\frac{ 18 }{ 11 }\]

OpenStudy (shamim):

\[so x=0,\frac{ 18 }{ 11 }\]

OpenStudy (shamim):

r u there

OpenStudy (mew55):

yes that is correct XD

OpenStudy (shamim):

ok

OpenStudy (anonymous):

thanks

OpenStudy (shamim):

u r most welcome